OpenStudy (anonymous):

What will the last payment be to pay off a $1,250 computer plus 8% sales tax with monthly payments of $120?   $10 $11 $12 $20 $30

OpenStudy (anonymous):

1250*8%=100 1250+100=1350 1350=11*120+30 so you will pay 120$ for 11 months and the last payment will be 30$.